OpenStudy (anonymous):

10x-99>(10/x) to solve the inequality, is it (-infin,-1/10)u(10,infin)?

OpenStudy (anonymous):

\[10x-99>\frac{10}{x}\\ 10x^2-99x-10>0\\ (10x+1)(x-10)>0\] What does this tell you?

OpenStudy (anonymous):

and is 0 included in the set, or no? I say no.

OpenStudy (anonymous):

-1/10<x<0 and x>10

OpenStudy (anonymous):

but the notation i think is confusing me.

OpenStudy (anonymous):

because i know (10,infin) but i don't know how to express the -1/10

OpenStudy (anonymous):

Yep, that's the right answer! \(x>10\) is the same as \((10,\infty)\), like you said. As for \(-\frac{1}{10}<x<0,\) the interval would be written similarly: \((-\frac{1}{10},0),\) since neither endpoint is included. The answer would then be the union of the two intervals.

OpenStudy (anonymous):

yayyyy. thank you! :].

OpenStudy (anonymous):

You're welcome